# Review-based Recommender Systems: A Survey of Approaches, Challenges and Future Perspectives

### v2

@@ Abstract.
Recommender systems play a pivotal role in helping users navigate an overwhelming selection of products and services. On online platforms, users have the opportunity to share feedback in various modes, including numerical ratings, textual reviews, and likes/dislikes. Traditional recommendation systems rely on users’ explicit ratings or implicit interactions (e.g., likes, clicks, shares, saves) to learn user preferences and item characteristics. Beyond these numerical ratings, textual reviews provide insights into users’ fine-grained preferences and item features. Analyzing these reviews is crucial for enhancing the performance and interpretability of personalized recommendation results. In recent years, review-based recommender systems have emerged as a significant sub-field in this domain. In this paper, we provide a comprehensive overview of the developments in review-based recommender systems over recent years, highlighting the importance of reviews in recommender systems, as well as the challenges associated with extracting features from reviews and integrating them into ratings. Specifically, we present a categorization of these systems and summarize the state-of-the-art methods, analyzing their unique features, effectiveness, and limitations. Finally, we propose potential directions for future research, including the integration of multimodal data, multi-criteria rating information, and ethical considerations.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: POMACSjournalvolume: XXjournalnumber: Xarticle: XXXpublicationmonth: 4ccs: Recommender Systemccs: Customer review
@@ 1. Introduction
In the rapidly advancing digital era, the sheer volume of online content, including products, movies, restaurants, and services, presents a significant challenge: information overload. This overwhelming amount of information often makes it difficult for people to find what they truly need or prefer. In response, recommender systems or information filtering techniques have emerged as key solutions, tackling the issue of consumer choice by guiding users to content that aligns with their interests and preferences. These systems play an important role in providing personalized recommendations, enhancing user experiences, contributing to business growth, and simplifying the decision-making process. These systems are extensively implemented by major businesses such as Google, Facebook, and Amazon, and across a wide range of web domains, including online shopping platforms (Schafer et al., 1999; Hwangbo et al., 2018; Zhu et al., 2014), healthcare services (Tran et al., 2021), educational resources (Tan et al., 2008), news portals (Raza and Ding, 2019a), and media streaming services (Majid et al., 2013).
@@
Collaborative Filtering (CF) (Su and Khoshgoftaar, 2009) and Content-based Filtering (CBF) (Thorat et al., 2015) have been fundamental to developing personalized recommender systems. CF excels when ample user ratings are available but struggles with sparse data, leading to inaccuracies and a lack of transparency. Content-based approaches aim to mitigate some of CF’s limitations by representing users and items through various data forms, focusing on item attributes and user preferences. Despite these efforts, content-based approaches encounter their own challenges due to the dynamic nature of user interests and content. With the proliferation of review data in online platforms, a new avenue for recommendation systems has emerged. Unlike traditional recommender systems that primarily rely on numerical ratings or browsing history, review-based systems explore the rich textual feedback provided by users. Review-based systems (Wu et al., 2019a; Liu et al., 2019a; Zheng et al., 2017; Chen et al., 2015; Seo et al., 2017; Chin et al., 2018; Huang et al., 2020a; Al-Ghuribi and Noah, 2019) leverage the wealth of information contained within user reviews, offering a promising solution to the limitations of the traditional rating-based methods.
@@
Review-based recommender systems harness the power of user reviews, extracting fine-grained aspects and associated sentiments, to offer more personalized, accurate, and explainable recommendations (Wu et al., 2019a). However, incorporating user reviews into recommendation models brings unique challenges. These include extracting fine-grained user preferences, and item characteristics from reviews, and leveraging them to learn the overall preferences. In addition, handling the vast and varied nature of textual data, and addressing issues like bias and privacy concerns are important (Raza et al., 2024). Hence, review-based recommender systems require advanced methodologies and innovative solutions.
Over the past years, extensive research has been conducted on recommender systems, including various systematic reviews (Adomavicius and Tuzhilin, 2005; Konstan and Riedl, 2012; Burke, 2002; Shi et al., 2014; Jannach et al., 2015; Khan et al., 2017; Batmaz et al., 2019). However, there appears to be a gap in comprehensive reviews specifically targeting review-based recommender systems. Previous surveys have broadly covered traditional recommender systems, deep learning based systems (Zhang et al., 2019), recommender systems using graph neural networks (Gao et al., 2023; Wu et al., 2022), multi-criteria recommender systems (Al-Ghuribi and Noah, 2019), session-based recommender systems (Wang et al., 2021a) and more recently reinforcement learning-based recommendation systems (Afsar et al., 2022), but have not concentrated exclusively on review-based systems. Bobadilla et al. (Bobadilla et al., 2013) presents an extensive overview of recommender systems and collaborative filtering, discussing their evolution and future directions. Lu et al. (Lu et al., 2015) surveyed application developments in recommender systems, categorizing applications and summarizing key algorithms. There are also survey papers that focus on context-aware recommendation systems, offering overviews of recent developments and future challenges (Haruna et al., 2017; Raza and Ding, 2019b).
A significant insight on deep learning-based recommender systems are presented in (Zhang et al., 2019), summarizing the state-of-the-art developments and current trends in the field. The study highlights the advantages and disadvantages of deep learning applications in this area. A comprehensive survey on deep learning focused knowledge graph is presented in (Gao et al., 2020a). The paper exclusively emphasizes on state-of-the-art frameworks (e.g. graph embeddings), and addresses the challenges such scalability, and cold-start issues (Gao et al., 2020a). This survey also summarizes commonly used datasets, evaluation methods, and open-source codes, while proposing potential research directions. More recently, the latest developments and trends in recommendation systems have been explored to include the consideration of robustness, data bias, and fairness, alongside with a brief review of personalized and group recommender systems (Li et al., 2023a)
Despite these extensive surveys, the specific area of review-based recommender systems is overlooked. This paper aims to fill this gap by focusing exclusively on review-based recommender systems. Here, we explore various methods for feature extraction from reviews, categorize advanced techniques, identify key challenges and solutions specific to this domain. We also discuss future possibilities in review-based recommender systems, offering insights into their development, methodologies, and potential advancements. Thus, our survey, with its focused approach, stands out from previous general surveys by providing in-depth insights into the rapidly evolving field of review-based recommender systems.
Contribution of this survey: The key contributions of this survey can be summarized as follows:
@@
    We systematically explore review-based recommender systems and introduce a classification scheme that considers feature extraction techniques and methodologies used to incorporate reviews into the recommendation process. To the best of our knowledge, this is the first comprehensive survey in this particular domain.
@@
    We provide a detailed overview of the state-of-the-art models, analyzing their unique features, effectiveness, limitations, and evaluation methods.
@@
    We investigate significant challenges and recent developments, proposing innovative directions for future research that address practical and theoretical gaps in the field. Our findings suggest pathways for developing more robust and transparent systems.
@@
Methodology: To conduct a systematic review on review-based recommender systems, we collect the relevant papers from a wide range of venues including ACM conference proceedings (e.g., RecSys, SIGIR, CIKM, AAAI, WWW, WSDM, KDD, UMAP, EMNLP, ACL), Google Scholar, and IEEE Xplore. Our review period spans from January 2015 to January 2024. We use specific search keywords that include ”review-based recommender systems”, ”review-aware recommender systems”, ”aspect-based recommender systems”, and ”aspect-aware recommender systems”. For conference proceedings, we initially select all papers that contain any of the above mentioned keywords in their titles, resulting in a total of 14,862 papers obtained across all conferences. In contrast, for Google Scholar and IEEE Xplore, we Utilize different strategy e.g. Boolean Operators. Our search queries are structured as (”review-based” OR ”review-aware” OR ”aspect-based” OR ”aspect-aware”) AND ”recommender systems”. Initial search results reveal a total of 5140 papers.
In contrast, IEEE Xplore reports much lower counts for similar keywords, e.g 855 papers. We perform a preliminary screening based on titles and abstracts, focusing on relevance to our study’s goals and careful selection. This screening involves two reviewers to enhance the reliability of the selection process, with discrepancies resolved through discussion, ultimately shortlisting 71 papers for detailed analysis.
@@
These selected papers undergo a thorough content analysis to extract information on methodologies, findings, and trends’ evolution over time, using both qualitative and quantitative techniques to ensure a robust interpretation of the literature. Figure 1 illustrates the distribution of the selected publications across different sources and formats, as well as their annual distribution, helping to understand the focus and growth of research areas within the field. Considering the rapid advancements in review-based recommender systems, we plan to update our literature review annually on arXiv, incorporating recent publications to stay updated with the latest developments.
@@
Figure 1. Trends in Review-Based Recommender Systems Research from 2015 to 2024
@@ 2. Background
Before further exploring the review-based recommender systems, we will discuss each of these categories in detail. Recommender systems are broadly classified into three categories: Collaborative Filtering (Schafer et al., 2007), Content-based Filtering (Pazzani and Billsus, 2007), and hybrid systems (Thorat et al., 2015). In the following subsections, we briefly discuss each of the categories.
@@ 2.1. Collaborative Filtering
Collaborative Filtering (CF) (Schafer et al., 2007) is a method used in recommender systems to predict a user’s interests based on the preferences of a larger user group. This technique relies on collecting and analyzing data on user behaviors, activities, or preferences to make personalized recommendations (Su and Khoshgoftaar, 2009). The core idea behind collaborative filtering is that, if two users have demonstrated similar preferences in their interactions with items in the past, it is likely that their future interaction patterns will also be similar. CF establishes a relationship between two different kinds of entities: users and items.
CF can be categorized into two main types: Memory-based (Deshpande and Karypis, 2004; Pu et al., 2011), and Model-based approaches (Paterek, 2007; Sarwar et al., 2001). Memory-based approaches make predictions about a user’s interests based on the similarity between users or items. This involves directly using the raw data of past user interactions to predict future preferences. On the other hand, model-based approaches involve building and training a model using user-item interaction data. These models, once trained, can predict a user’s interest in an item based on the learned patterns. The prediction in CF systems is primarily based on the historical data of users and their interactions with items. In memory-based methods, predictions are made by averaging the ratings of similar items or by considering the ratings of similar users. Model-based approaches, however, predict user preferences by learning the latent factors of users and items from the interaction data, often employing techniques like matrix factorization (Kalman, 1996).
Recently, the trend in CF has shifted towards more sophisticated model-based approaches, incorporating advanced machine learning and deep learning models. Neural Collaborative Filtering (NCF) models (Chen et al., 2018a), for instance, use neural networks to learn the complex non-linear relationships between users and items.

### v3

@@ Abstract.
Recommender systems play a pivotal role in helping users navigate a vast selection of products and services. On online platforms, users have the opportunity to share feedback in various modes, such as numerical ratings, textual reviews, and likes/dislikes. Traditional recommendation systems rely on users’ explicit ratings or implicit interactions (e.g., likes, clicks, shares, and saves) to learn user preferences and item characteristics. Beyond numerical ratings, textual reviews provide insights into users’ fine-grained preferences and item features. Analyzing these reviews is crucial for enhancing the performance and explainability of personalized recommendation results. In this paper, we provide a comprehensive overview of the development in review-based recommender systems over recent years, highlighting the importance of reviews in recommender systems, as well as the challenges associated with extracting features from reviews and integrating them into ratings. Specifically, we introduce a classification scheme in terms of both the integration of reviews into recommendation systems and the technical methodology. Additionally, we summarize the state-of-the-art methods, analyzing their unique features, effectiveness, and limitations. The study also presents the various evaluation metrics, comparative analysis, datasets, and real-world applications of review-based recommendation systems. Finally, we propose potential directions for future research, including multi-modal data integration, multi-criteria rating information, and ethical considerations.
@@
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: POMACSjournalvolume: XXjournalnumber: Xarticle: XXXpublicationmonth: 5ccs: Recommender Systemccs: Customer review
@@ 1. Introduction
In the rapidly advancing digital era, the sheer volume of online content, including products, movies, restaurants, and services, presents a significant challenge: information overload. This overwhelming amount of information often makes it difficult for people to find what they truly need or prefer. In response, recommender systems have emerged as key solutions, guiding users to content that aligns with their interests and preferences. These systems play an important role in providing personalized recommendations, enhancing user experiences, contributing to business growth, and facilitating the decision-making process (Huang et al., 2020a). Big technology companies including Google, Meta, Netflix and Amazon, extensively use these systems across a wide range of web domains such as e-commerce (Schafer et al., 1999; Hwangbo et al., 2018), healthcare (Tran et al., 2021), education (Tan et al., 2008), news portals (Raza and Ding, 2019a), and media streaming (Majid et al., 2013).
@@
Traditional approaches to recommender systems such as Collaborative Filtering (CF) (Su and Khoshgoftaar, 2009) and Content-based Filtering (CBF) (Thorat et al., 2015) have been used for developing personalized recommender systems. CF performs well when there is ample user rating data, but it struggles with sparse data, leading to inaccuracies and a lack of transparency (Xu et al., 2018). CBF addresses some limitations of CF by representing users and items through item attributes and user preferences (Mooney and Roy, 2000), but they also face challenges due to the dynamic nature of user interests (Thorat et al., 2015).
The rise of user reviews on online platforms has introduced a new dimension to recommendation systems. Unlike traditional systems that primarily rely on numerical ratings or browsing history, review-based systems leverage rich textual feedback, covering various elements such as product aspects, sentiment, and even writing style (Dang et al., 2021; Hernández-Rubio et al., 2019; Artemenko et al., 2020; Al-Ghuribi et al., 2023a; Hazar et al., 2022). Aspect-based opinion mining further enhances these systems by extracting detailed product attributes from reviews and aligning them with user preferences for more tailored recommendations (Bhattacharya et al., 2022; Hernández-Rubio et al., 2019). Additionally, sentiment analysis captures user attitudes and emotions, providing a deeper understanding of preferences (Dang et al., 2021; Bhattacharya et al., 2022; Musto et al., 2017). These systems incorporate factors such as review structure, content richness, and contextual information to address the limitations of rating-based methods (Wu et al., 2019a; Liu et al., 2019a; Zheng et al., 2017; Chen et al., 2015; Seo et al., 2017; Chin et al., 2018; Huang et al., 2020a; Al-Ghuribi and Noah, 2019; Osman et al., 2021). For example, in e-tourism, reviews can systematically be summarized for user preferences, leading to more accurate and targeted mobile recommendations (Artemenko et al., 2020).
@@
Review-based recommender systems incorporate user-generated reviews to offer more personalized, accurate, and explainable recommendations (Rokach et al., 2022; Wu et al., 2019a). However, incorporating textual data into recommendation models presents unique challenges, such as extracting fine-grained user preferences and item characteristics, handling diverse and abundant textual data, and addressing concerns around bias and privacy (Raza et al., 2024a). Furthermore, systems must now account for cross-domain applications and multi-language challenges, where multilingual reviews provide an additional complexity that recent models have begun to address using multilingual aspect-based sentiment analysis (Liu et al., 2021d). As a result, review-based recommender systems require advanced methodologies and innovative solutions to tackle these evolving challenges.
Recently, hybrid models that combine CF with sentiment analysis and aspect-based methods have emerged as key developments. These approaches aim to overcome limitations of the traditional models by incorporating both structured and unstructured data (Ziani et al., 2017; Hung, 2020; Osman and Noah, 2018; Mishra et al., 2020; Contratres et al., 2018; Asani et al., 2021). Utilizing natural language processing, and deep learning techniques into review-based systems has driven the development of sophisticated recommendation models (Shoja and Tabrizi, 2019; Preethi et al., 2017; Hu et al., 2020; Chiranjeevi and Rajaram, 2023; Sharma et al., 2024). Multi-criteria systems, which consider different aspects such as service, price, and quality, have shown to improve both recommendation accuracy and user satisfaction (Musto et al., 2017; Singh et al., 2024). Recent studies have demonstrated how aspect-based sentiment analysis, combined with CF, can provide more nuanced recommendations, particularly in large-scale datasets (Ozcan et al., 2022). Beyond aspects, elements such as sentiment polarity, opinion diversity, and even narrative structure contribute to a holistic understanding of user preferences.
While extensive research has reviewed traditional recommender systems (Li et al., 2023a; Lu et al., 2015; Adomavicius and Tuzhilin, 2005; Konstan and Riedl, 2012; Burke, 2002; Shi et al., 2014; Jannach et al., 2015; Khan et al., 2017; Batmaz et al., 2019), deep learning-based systems (Zhang et al., 2019b), graph-based approaches (Gao et al., 2023; Wu et al., 2022; Gao et al., 2020a), multi-criteria methods (Al-Ghuribi and Noah, 2019; Musto et al., 2017), session-based systems (Wang et al., 2021a; Huang et al., 2021), context-aware recommender systems (Haruna et al., 2017; Raza and Ding, 2019b), knowledge-based (Bobadilla et al., 2013) and reinforcement learning systems (Afsar et al., 2022; Wang et al., 2020), there remains a notable gap in surveys specifically focusing on review-based recommender systems. Although review-based recommender systems have been partially explored in earlier works (Chen et al., 2015; Al-Ghuribi and Noah, 2019), these studies are now quite outdated. For instance, Chen et al. (Chen et al., 2015) conducted a survey highlighting the importance of creating user and item profiles by leveraging user-generated reviews. However, this survey is limited to research works utilizing reviews for user profile construction and multi-faceted opinion mining. Additionally, being published in 2015, the survey does not reflect recent progress in deep learning and NLP techniques (e.g., attention mechanism (Tal et al., 2021), transformer models (Vaswani, 2017), LLMs (Li et al., 2023b), Contrastive Learning (Xia et al., 2022), Reinforcement Learning (Wang et al., 2020), Graph Contrastive Learning (You et al., 2020), Counterfactual arguments (Tan et al., 2021), etc.) to learn users’ fine-grained preferences and item attributes. Similarly, a survey on multi-criteria review-based recommendations was conducted by Al-Ghuribi and Noah (Al-Ghuribi and Noah, 2019) in 2019, leveraging user-generated reviews to improve the performance of multi-criteria recommendation systems. Although this survey explores fine-grained sentiment analysis, it ignores several critical aspects of review-based recommendation systems, including the integration techniques of reviews with ratings, sentiment polarity at different levels, comprehensive user and item feature learning, and advanced representation learning techniques. In 2021, Shalom et al. (Shalom et al., 2021) surveyed the application of NLP in recommendation systems. Their survey primarily focused on the use of reviews and explanation generation, while ignoring the detailed coverage of multifaceted representation methods and advanced feature learning techniques. Moreover, it does not focus on the latest advancements in review-based recommendation systems.
This paper bridges these gaps by offering a comprehensive review that introduces review integration strategies, analyzes the latest feature learning and deep representation learning methods, investigates domain-specific challenges, and suggests future research directions. By covering both aspect-based and broader review-aware techniques, this survey provides a holistic perspective on this rapidly evolving field (Margaris et al., 2020). The key contributions of this survey can be summarized as follows:
@@
    We systematically explore review-based recommender systems and introduce two classification schemes—one based on how reviews are incorporated into recommendation systems and the other on technical methodologies for learning review features. To the best of our knowledge, this is the first comprehensive survey focused on this domain.
@@
    We provide a comprehensive overview of the state-of-the-art models, analyzing unique features, effectiveness, limitations, datasets, evaluation methods, and real-world applications of recommendation systems.
@@
    We address key challenges and recent developments, proposing research directions to tackle theoretical and practical gaps, focusing on developing robust and transparent systems.
@@
Methodology: To conduct a systematic review on review-based recommender systems, we collected the relevant papers from a wide range of sources, including proceedings of the prestigious conferences in this field (e.g., RecSys, SIGIR, CIKM, AAAI, WWW, WSDM, KDD, UMAP, EMNLP, ACL), Google Scholar, and IEEE Xplore. Our review period spans January 2015 to October 2024. We used specific search keywords such as “review-based recommender systems”, “review-aware recommender systems”, “aspect-based recommender systems”, “aspect-aware recommender systems”, and “sentiment analysis in recommender systems”.
For conference proceedings, we initially recorded a total number of papers that contain any of the above-mentioned keywords in their titles, resulting in a total of 14,862 papers across all conferences. For Google Scholar and IEEE Xplore, we utilized a different strategy, applying Boolean operators. Our search queries were structured as ( “review-based” OR “review-aware” OR “aspect-based” OR “aspect-aware” OR “user reviews” OR “user review elements” OR “sentiment analysis” ) AND “recommender systems”. The initial Google Scholar search returned a total of 17,200 papers, whereas IEEE Xplore reported a significantly lower count of 855 papers for a similar keyword search. After the initial search, we skimmed through the titles and created a preliminary shortlist consisting of 529 papers. In the next step, we performed a screening based on titles and abstracts, focusing on relevance to the goals of our study and ensuring a careful selection process. This screening involved two reviewers to enhance reliability, with discrepancies resolved through discussion, ultimately shortlisting 113 papers for detailed analyses from all sources.
@@
The selected papers undergo a thorough content analysis to extract information on methodologies, findings, and the evolution of trends over time. Figure 1 illustrates the distribution of the selected publications across different sources and formats, as well as their annual distribution.
[…]
@@
Trends in Review-based Recommender Systems (2015-–2024): Key publications in journals and conferences
Figure 1. Trends in Review-based Recommender Systems (2015-–2024): Key publications in journals and conferences
To ensure the inclusion of the most relevant and impactful research, we applied the following inclusion criteria: (1) The paper must explicitly focus on review-based, review-aware, or aspect-based recommender systems. (2) The paper should introduce a novel methodology or present a significant enhancement to existing methods. (3) The paper must be peer-reviewed and published in a recognized NLP, AI or machine learning conference, such as those listed earlier. (4) Additionally, papers that explore the intersection of sentiment analysis and recommender systems, especially those utilizing user reviews to improve recommendation quality, were included. (5) For recent papers on arXiv, we required that they have at least 5 citations per year to ensure their influence in the field. (6) Lastly, we included several highly cited papers even if they were not from widely recognized venues.
For exclusion, we applied the following criteria: (1) Papers that focus solely on traditional collaborative filtering or content-based filtering without integrating user reviews or textual data were excluded. (2) Papers that dealt exclusively with non-recommender system applications, even if they involved sentiment analysis or aspect extraction, were also excluded. By following these criteria, we aimed to focus on research that was most relevant to the evolution of review-based recommender systems, ensuring that the selected papers would contribute meaningfully to our survey of the field.
@@ 2. Background
Recommender systems are broadly classified into three categories: Collaborative Filtering (Schafer et al., 2007), Content-based Filtering (Pazzani and Billsus, 2007), and Hybrid Systems (Thorat et al., 2015). Before further exploring the review-based recommender systems, we will discuss each of these categories in detail in the following subsections.
@@ 2.1. Collaborative Filtering
Collaborative Filtering (CF) (Schafer et al., 2007) is a method used in recommender systems to predict a user’s interests based on the preferences of a larger user group. This technique relies on collecting and analyzing data on user behaviors, activities, or preferences to make personalized recommendations (Su and Khoshgoftaar, 2009). The core idea behind CF is that, if two users have shown similar preferences in their past interactions with items, they are likely to continue showing similar patterns in the future. CF establishes a relationship between two entities: users and items.
CF can be categorized into two main types: Memory-based (Deshpande and Karypis, 2004; Pu et al., 2011), and Model-based approaches (Paterek, 2007; Sarwar et al., 2001). Memory-based approaches make predictions based on similarities between users or items, directly using past user interaction dat. Model-based approaches, on the other hand, involve building and training models from user-item interaction data to predict a user’s interest in an item based on learned patterns. In CF systems, prediction is primarily based on historical data of user-item interactions. Memory-based methods make predictions by averaging the ratings of similar items or users, while model-based approaches predict user preferences by learning latent factors of users and items, often using techniques like matrix factorization (Kalman, 1996).
Matrix factorization maps both the items and users by vectors of factors inferred from item ratings. To illustrate the matrix factorization technique, consider a user $u$ with associated vector $\mathbf{p}_{u}\in\mathbb{R}^{d}$ and an item $i$ with associated vector $\mathbf{q}_{i}\in\mathbb{R}^{d}$, where $d$ is the dimension of the vector. The dot product of these vectors represents the interaction between user and item, estimating the user’s overall preference toward the item, as follows:
(1) $\hat{r}_{ui}=\mathbf{p}_{u}^{T}\mathbf{q}_{i}+b$
where b is the bias term. The core idea behind matrix factorization is to minimize the difference between the predicted rating score $\hat{r}_{ui}$ and the true rating score $r_{ui}$. This is done by solving the following objective function:
(2) $\min_{P,Q}\sum_{(u,i)\in R}(r_{ui}-\mathbf{p}_{u}^{T}\mathbf{q}_{i})^{2}+\lambda(\|P\|^{2}+\|Q\|^{2})$
where $P$ is an $m\times k$ matrix representing user features with $k$ number of latent factors, $Q$ is a $k\times n$ matrix representing item features, and $\lambda\left(\|P\|^{2}+\|Q\|^{2}\right)$ is the regularization term with $\lambda$ representing regularization parameter. The ultimate objective of the matrix factorization is to reconstruct the original matrix $R$ by the product of two smaller matrices $P\times Q$ as closely as possible.